Hi i need help on solving using substitution for two problems:
x+y=1
2x+3y=-4
Please show all work thank you!

Answer by rfer(16322)

y=-x+1
2x+3y=-4
2x+3(-x+1)=-4
2x-3x+3=-4
-x=-7
x=7
-----------
2(7)+3y=-4
14+3y=-4
3y=-4-14
3y=-18
y=-18/3
y=-6
(7,-6)
-------------
x+y=1
7-6=1